OpenStudy (anonymous):

Find all values of t in the interval [0, 2pi] that satisfy the given equation: sin t=square root 2/2

OpenStudy (campbell_st):

so then \[x = \sin^{-1}(\frac{2\sqrt{2}}{2}) = \frac{\pi}{4}\] so you have the 1st quadrant angle, you just need to find the 2nd quadrant angle which requires \[\pi - x\]